For the plaque assay, your T4 bacteriophage titer is 7.8x108 pfu/mL. You do six 1:10 dilutions of the virus, putting 10µL of each dilution in a tube of top agar with 100µL of E. coli. After pouring the contents of the top agar onto an LA plate, you incubate overnight. How many plaques would you expect to see on the final plate?
Answer #1

Plaque forming unit/ml = number of plaque / dilution * volume plated.

Plaque forming unit/ml =  7.8x108

six 1:10 dilutions were perfromed. So overall dilution is 1/10*1/10*1/10* 1/10*1/10*1/10 = 1/1000000 = 10-6

Volume plated = 10ul = 0.01ml.

  7.8x108 = number of plaque /10-6 *0.01

=Number of plaque =  7.8x108 * 10-6 *10-2

= 7.8x108 * 10-8

= 7.8.

So we would expect to see 7 - 8 plaque on the final plate.

Please write to me if you need more help.